What is the 9th term of the following geometric sequence? Enter a whole number.
7/9,-7/3, 7, -21, 63, ...

Answer by richwmiller(17219)

(n-1) term is multiplied by -3 to get the nth term
-3*63=-189 is the next term
T9 = 7/9 * (-3)^(8)=5103
Tn = T1 * r^(n - 1)